OpenStudy (anonymous):

What is the solution of the linear-quadratic system of equations? y= x^2 + 5x - 3 y-x =2

OpenStudy (anonymous):

I am totally not getting this...help!

OpenStudy (anonymous):

Solve one of the equations for either y or x and then plug that into the other equation for the respective variable

OpenStudy (anonymous):

So to get y by itself, would I change the bottom equation to y = x +2?

OpenStudy (anonymous):

y-x=2 y=x+2 x^2+5x-3=x+2

OpenStudy (anonymous):

now try to factorise

OpenStudy (anonymous):

y = x - 3?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry Im not good at factorising.

OpenStudy (anonymous):

X^2+5x-3=x+2 X^2+5x-x-3-2=0 got it?

OpenStudy (anonymous):

x^2+4x-5=0

OpenStudy (anonymous):

now see, 5+-1=4 -1*5=-5

OpenStudy (anonymous):

so x^2+4x-5=(x+5)(x-1)=0

OpenStudy (anonymous):

now can u find values of x?

OpenStudy (anonymous):

so now I substitute to find x? Im sorry, I have never been good at this.

OpenStudy (anonymous):

(x+5)(x-1)=0. understood this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and if x-1 = 0 then x =1

OpenStudy (anonymous):

of course. now we have 2 values of x. can you find y with that?

OpenStudy (anonymous):

Yeah I think so. I'll go try it

OpenStudy (anonymous):

x=-5 or x=1

OpenStudy (anonymous):

that's nice, u r a gud student, i feel. bye

OpenStudy (anonymous):

thanks. So x = 7 or 3?

OpenStudy (anonymous):

I mean Y = 7 or 3?

OpenStudy (anonymous):

for x=-5, y--5=2 y+5=2 y=2-5 other value is right.

OpenStudy (anonymous):

- x*- x= +x^2

OpenStudy (anonymous):

x--x=x+x. got it?